How many moles of H2CO3 are in a 1.44 kg sample of the compound?

A. 23.22 moles
B. 43.07 moles
C. 0.043 moles
D. 0.23 moles
E. None of the Above
1
Expert's answer
2018-01-17T05:45:50-0500
m(H2CO3) = 1.44 kg = 1440 g
M(H2CO3) = 1*2+12+16*3 =62 g/mol
n(H2CO3) = m(H2CO3)/M(H2CO3)
n(H2CO3) = 1440 g / 62 g/mol = 23.22 moles
A. 23.22 moles